I keep getting the answer wrong please show steps!
Part A The change in enthalpy (AH) for a reaction is -23.5 kJ/mol. The equilibrium constant for the reaction is 2.1x103 at 298 K What is the equilibrium constant for the reaction at 605 K? Express your answer using two significant figures. Submit Request Answer
Answer #1

ln (K2/K1) = dH/R (1/T1 - 1/T2)

K1 = 2.1 x 103 ,T1 = 298 K

K2 = ? , T2 = 605 K

dH = -23500 J/mol , R = 8.314 J/mol.K

ln (K2/2100) =(-23500 J/mol) /(8.314 J/mol.K) (1/298 - 1/605)1/K

ln K2 - 7.65 = -2826.56 (0.00170)

ln K2 = 2.8369

K2 = 17.06